Marion is 5 years younger than Susan. In 10 years Susan will be 3 times as old as Marion was 3 years ago. Find their present ages. The answer is Susan is 17 and Marion is 12. I just need the solution to this.
Answer by macston(5194)

Susan's age = S Marion's age=M=S-5
S+10=3(M-3) Substitute above equation for M
S+10=3((S-5)-3)
S+10=3(S-8)
S+10=3S-24 Subtract S from each side
S+10-S=3S-24-S
10=2S-24 Add 24 to each side
10+24=2S-24+24
34=2S divide each side by 2
17=S ANSWER 1 Susan is 17 years
M=S-5=17-5=12 years ANSWER 2 Marion is 12 years
CHECK
In 10 years, Susan will be 3 times as old as Marion was 3 years ago
In 10 years, Susan will be 27. 3Years ago Marion was 9
3*9=27
27=27